You buy items costing $800 and finance the cost with a fixed loan installment for 18 months at 5% simple interest per year. What is the finance charge? What is the monthly payment?


1
Expert's answer
2021-07-13T15:21:22-0400

"A = P\times(1+rt)"

Where:

A = Total amount

P = principal amount 

r = interest rate

t = time period

 we will first divide the interest rate by 12 to find the monthly interest rate.

"r= \frac{5\%}{12} = 0.41667\%"

Now we will find the total amount (interest + principal)

"A = P\times(1+rt)\\

= 800 \times (1+(0.0041667\times18))\\

= 800 \times(1+ 0.075)\\

= 800 \times 1.075\\

= 860"

Thus,

Total amount payable with interest is $860.

To find the finance charges, we will simply subtract the principal amount from the total amount payable found in previous step.

Finance charges

"= \$860 - \$800

= \$60."

Thus,

Finance charges are $60. 

Now to find the monthly payment, we will simply divide the total amount payable by number of months.

Monthly payment

"= \frac{860}{ 18}\\

= 47.78"

= $48 (rounded)

Thus,

Monthly payment on the loan will be $48.